# Computing Exact Shapley Values in Polynomial Time for Product-Kernel Methods

### v1

@@ 3.1 Functional decomposition and game-theoretic analysis
A decision function constructed using a product kernel naturally admits a functional decomposition where each component $f_{\mathcal{S}}(\boldsymbol{x}_{\mathcal{S}})$ is given by:
$f_{\mathcal{S}}(\boldsymbol{x}_{\mathcal{S}})=\sum_{i=1}^{n}\alpha_{i}\prod_{j\in\mathcal{S}}\big{(}k_{j}(x_{j},x^{(i)}_{j})-1\big{)}.$ (3)
This theorem demonstrates that product-kernel methods induce a functional decomposition with a specific structure. It clarifies how different features contribute to the prediction and interact with each other. Leveraging this functional decomposition, we define $v_{\boldsymbol{x}}(\mathcal{S})$ as the aggregate of all components $f_{\mathcal{T}}(\boldsymbol{x}_{\mathcal{T}})$, $\forall\mathcal{T}\subseteq\mathcal{S}$, which quantifies the cumulative influence of the coalition $\mathcal{S}$ [23, 31].
Definition 2.
For a decomposable function $f$, the value function is $v_{\boldsymbol{x}}(\mathcal{S})=\sum_{\mathcal{T}\subseteq\mathcal{S}}f_{\mathcal{T}}(\boldsymbol{x}_{\mathcal{T}})$.
We now show that this value function can be efficiently computed for product-kernel learning methods.
Proposition 3.
For product-kernel learning methods, $v_{\boldsymbol{x}}(\mathcal{S})$ can be computed efficiently as
$v_{\boldsymbol{x}}(\mathcal{S})=\boldsymbol{\alpha}^{\top}k_{\mathcal{S}}(\mathbf{X}_{\mathcal{S}},\boldsymbol{x}_{\mathcal{S}}).$ (4)
The value function (4) quantifies the contribution of a feature subset $\mathcal{S}$ by evaluating the kernel function restricted to that subset only. Thus, it isolates the effect of $\mathcal{S}$ by applying the kernel only to those features, ensuring that the contribution of $\mathcal{S}$ is measured independently of the remaining features. Also, computing the value function has linear time complexity, making it highly efficient.
3.2 A recursive formulation for Shapley value
Next, we show that for the value function (4), the Shapley value for product-kernel learning methods can be recursively computed in polynomial time.
Theorem 4.
@@ 3.2 A recursive formulation for Shapley value
Let $\mathcal{Z}:=\{\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{d}\}$ with $\boldsymbol{z}_{i}=k_{i}(\mathbf{X}_{i},x_{i})$ and $e_{q}(\mathcal{Z})$ the elementary symmetric polynomials (ESPs) of order $q$ over $\mathcal{Z}$, defined recursively as $e_{q}(\mathcal{Z}_{-j})=\frac{1}{q}\sum_{r=1}^{q}(-1)^{r-1}e_{q-r}(\mathcal{Z}_{-j})\odot p_{r}(\mathcal{Z}_{-j})$, where $p_{r}(\mathcal{Z})=\sum_{\boldsymbol{z}_{i}\in\mathcal{Z}}\boldsymbol{z}_{i}^{r}$ is the element-wise degree-$r$ power sum. For product-kernel learning methods with coefficients $\boldsymbol{\alpha}$, the Shapley value $\phi_{j}^{\boldsymbol{x}}$ for feature $j$ of instance $\boldsymbol{x}$ can then be expressed as
@@
$\phi_{j}^{\boldsymbol{x}}:=\sum_{\mathcal{S}\subseteq\mathcal{D}\setminus\{j\}}\mu(|\mathcal{S}|)\big{(}v_{\boldsymbol{x}}(\mathcal{S}\cup\{j\})-v_{\boldsymbol{x}}(\mathcal{S})\big{)}=\boldsymbol{\alpha}^{\top}\left(\bigg{(}k_{j}(\mathbf{X}_{j},x_{j})-\boldsymbol{1}\bigg{)}\bigodot\sum_{q=0}^{d-1}\mu(q)e_{q}\big{(}\mathcal{Z}_{-j}\big{)}\right).$ (5)
@@
The recursion follows from Newton’s identities for ESPs [34], which we explain further in Appendix B. The intuition to arrive at equation (5) is to use the multiplicative structure of the product kernel to factorize the marginal contribution $v_{\boldsymbol{x}}(\mathcal{S}\cup\{j\})-v_{\boldsymbol{x}}(\mathcal{S})$ as $\boldsymbol{\alpha}^{\top}((k_{j}(\mathbf{X}_{j},x_{j})-1)\odot k_{S}(\mathbf{X}_{\mathcal{S}},\boldsymbol{x}_{\mathcal{S}}))$. This structure allows us then to push the summation inside the inner product between the coefficients $\boldsymbol{\alpha}$ and the kernel evaluations, and express the total sum over all subsets $\mathcal{S}\subseteq\mathcal{D}\setminus\{j\}$ in terms of weighted ESPs $e_{q}(\mathcal{Z}_{-j})$, which can then be computed recursively in $O(d^{2})$. A similar trick has been used in the context of additive Gaussian processes [35], but to our knowledge, it has not been previously leveraged for computing Shapley values in polynomial time. We present the algorithm in Appendix D, along with a modification that is numerically stable.
@@
Next, we show the additivity of the explanation for the value function in equation (4) (see Appendix E for discussion on the additivity of explanations for different values for the null game).
@@
4 Explaining kernel-based statistical discrepancies
This section demonstrates the interpretation of the two kernel-based statistical discrepancies, HSIC and MMD, with product kernels. We discuss the additivity of the explanations in Appendix E.
4.1 Distributing the discrepancy: Explaining MMD
[…]
$\widehat{\text{MMD}}^{2}(\mathbb{P},\mathbb{Q})=\frac{1}{n(n-1)}\sum_{i\neq j}k(\boldsymbol{x}^{(i)},\boldsymbol{x}^{(j)})+\frac{1}{m(m-1)}\sum_{i\neq j}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})-\frac{2}{nm}\sum_{i,j}k(\boldsymbol{x}^{(i)},\boldsymbol{z}^{(j)}).$
In this section, we propose an attribution method to allocate the overall discrepancy measured by MMD among the involved variables. Such an attribution is useful in many problems, including explaining MMD-based statistics for hypothesis testing (e.g., Gretton et al. [9]) and determining the contribution of variables to covariance shift (e.g., Zhang et al. [36]). By analyzing MMD through a game-theoretic perspective, we uncover its functional decomposition structure, which facilitates the formulation of a cooperative game under the assumption that $k$ is a product kernel.
Theorem 6.
The MMD with a product kernel has the following properties:
  • (i)
    The MMD function can be decomposed into contributions from all feature subsets $\mathcal{S}\subseteq\mathcal{D}$ as
    $\widehat{\text{MMD}}^{2}(\mathbb{P},\mathbb{Q})=\sum_{\mathcal{S}\subseteq\mathcal{D}}\bigg{[}\frac{1}{n(n-1)}\sum_{i\neq j}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x^{(i)}_{q},x^{(j)}_{q})-1\big{)}$
    $\qquad+\frac{1}{m(m-1)}\sum_{i\neq j}\prod_{q\in\mathcal{S}}\big{(}k_{q}(z^{(i)}_{q},z^{(j)}_{q})-1\big{)}-\frac{2}{nm}\sum_{i,j}\prod_{q\in\mathcal{S}}\big{(}k_{q}(x^{(i)}_{q},z^{(j)}_{q})-1\big{)}\bigg{]}.$
  • (ii)
    Given this decomposition, the corresponding value function, which represents the contribution of the variables in $\mathcal{S}$, is uniquely defined as follows: $v_{\text{MMD}}(\mathcal{S})=\frac{1}{n(n-1)}\sum_{i\neq j}k_{\mathcal{S}}(\boldsymbol{x}^{(i)}_{\mathcal{S}},\boldsymbol{x}^{(j)}_{\mathcal{S}})+\frac{1}{m(m-1)}\sum_{i\neq j}k_{\mathcal{S}}(\boldsymbol{z}^{(i)}_{\mathcal{S}},\boldsymbol{z}^{(j)}_{\mathcal{S}})-\frac{2}{nm}\sum_{i,j}k_{\mathcal{S}}(\boldsymbol{x}^{(i)}_{\mathcal{S}},\boldsymbol{z}^{(j)}_{\mathcal{S}}).$
@@ 4.1 Distributing the discrepancy: Explaining MMD
4.2 Distributing the dependence: Explaining HSIC
The Hilbert-Schmidt Independence Criterion (HSIC) is a kernel-based dependence measure that quantifies the statistical dependence between two random variables. Let $X$ and $Y$ be two random variables with $k(\cdot,\cdot)$ and $l(\cdot,\cdot)$ as reproducing (product) kernels defined on them. Then, $\text{HSIC}(X,Y):=\|\mathcal{C}_{XY}\|^{2}_{\text{HS}}$, where $\mathcal{C}_{XY}$ is the cross-covariance operator and $\|\cdot\|_{\text{HS}}$ is the Hilbert-Schmidt (HS) norm; see, e.g., [21, Sec. 3.6] for technical details. Given a sample $\{(\boldsymbol{x}^{(i)},\boldsymbol{y}^{(i)})\}_{i=1}^{n}\sim\mathbb{P}(X,Y)$, $\text{HSIC}(X,Y)$ can be estimated as $\widehat{\text{HSIC}}(X,Y)=(n-1)^{-2}\mathrm{tr}(\mathbf{K}\mathbf{H}\mathbf{L}\mathbf{H})$ where $\mathbf{K}\in\mathbb{R}^{n\times n}$ is the kernel matrix computed using the kernel $k$, i.e, $\mathbf{K}_{ij}=k(\boldsymbol{x}^{(i)},\boldsymbol{x}^{(j)})$, $\mathbf{L}\in\mathbb{R}^{n\times n}$ is the kernel matrix computed using the kernel $l$, i.e, $\mathbf{L}_{ij}=l(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})$, and $\mathbf{H}=I-\frac{1}{n}\boldsymbol{1}\boldsymbol{1}^{\top}$ is the centering matrix ensuring zero mean in the feature space.
Computing $\text{HSIC}(X,Y)$ gives us the overall dependence between $X$ and $Y$. We now focus on scenarios where $X$ represents a random variable of the feature vector $\boldsymbol{x}$ and $Y$ represents the scalar prediction outcome $y$. Our interest is then to distribute the overall dependence over the features in $\boldsymbol{x}$. This is particularly useful for feature selection and global sensitivity analysis, where the target is usually univariate.
The following theorem establishes that HSIC naturally induces a functional decomposition structure, enabling the formulation of a cooperative game.
Theorem 8.
For the HSIC with the product kernel on $X$, it has the following properties:
  • (i)
    The HSIC function can be decomposed into contributions from all feature subsets $\mathcal{S}\subseteq\mathcal{D}$ as
    $\widehat{\text{HSIC}}(X,Y)=\frac{1}{(n-1)^{2}}\sum_{\mathcal{S}\subseteq\mathcal{D}}\mathrm{tr}\bigg{(}\mathbf{H}\mathbf{L}\mathbf{H}\bigodot_{j\in\mathcal{S}}(\mathbf{K}_{j}-\boldsymbol{1}\boldsymbol{1}^{\top})\bigg{)}.$
  • (ii)
    Given this decomposition, the associated value function that quantifies the contribution of the variable set $\mathcal{S}$ to the overall dependency is uniquely defined as $v_{\text{HSIC}}(\mathcal{S})={(n-1)^{-2}}\mathrm{tr}\big{(}\mathbf{H}\mathbf{L}\mathbf{H}\mathbf{K}_{\mathcal{S}}\big{)}.$
Applying the same trick as in Proposition 7 to the value function $v_{\text{HSIC}}$ yields a similar recursive formulation of Shapley values for HSIC.
@@ 4.2 Distributing the dependence: Explaining HSIC
$\phi_{j}^{\text{HSIC}}=\frac{1}{(n-1)^{2}}\mathrm{tr}\left(\mathbf{H}\mathbf{L}\mathbf{H}\bigg{(}(\mathbf{K}_{j}-\boldsymbol{1}\boldsymbol{1}^{\top})\bigodot\sum_{q=0}^{d-1}\mu(q)E_{q}(\mathcal{K}_{-j})\bigg{)}\right).$
@@
The Shapley values $\phi_{j}^{\text{HSIC}}$ allow us to allocate the overall dependence between $X$ and $Y$ across individual features fairly, identifying the most influential ones for prediction. Our results can be generalized to scenarios when both $X$ and $Y$ are multivariate; see Appendix F for further details. The attribution of overall dependence to the involved multivariate variables has other applications in statistical inference as well, e.g., kernel-based (conditional) independence testing [14, 37].
5 Experiments
@@
We evaluate the effectiveness of PKeX-Shapley for product-kernel methods through a series of experiments conducted on a 24-core machine with 16GB RAM and an RTX4000 GPU. We further provide the experimental setups, training procedure, and extra experiments in Appendix G.
5.1 Demonstrating the effectiveness of recursion and the value function via local explanations
@@
Experiment 1: Recursion vs. Regression Formulation. We empirically demonstrate the advantage of PKeX-Shapley, compared to sampling-based approximations such as Kernel SHAP. We generate four synthetic datasets, each with $1000$ samples and $d\in\{10,20,30,50\}$ features. Features are independently drawn from a standard normal distribution, and target values are generated using a linear model with additive Gaussian noise ($\sigma=0.1$). For each dataset, we train a support vector regression model with an RBF kernel and use the trained model to compare explanation methods.
@@
For a fair comparison between the recursive and regression-based methods, we adopt the same value function as defined in Definition 2. Specifically, we first compute the exact Shapley values $\phi_{j}^{\boldsymbol{x}}$ using PKeX-Shapley, and then estimate the Shapley values using a regression-based approach analogous to Kernel SHAP, but modified to employ our value function from Definition 2. The estimator uses the paired coalition sampling scheme of Covert and Lee [38], with sample sizes ranging from $200$ to $10^{5}$, and computes $\hat{\phi}_{j}^{\boldsymbol{x}}$ by solving the corresponding weighted linear regression problem. To quantify the approximation error introduced by the regression formulation, we report the relative deviation error across $100$ randomly selected instances, defined as $\sum_{j=1}^{d}|\phi_{j}^{\boldsymbol{x}}-\hat{\phi}_{j}^{\boldsymbol{x}}|/|\phi_{j}^{\boldsymbol{x}}|$.
@@
Figure 1 plots the relative error as a function of sample size for each dataset over the 100 selected instances. When $d=10$, the relative error is already around 0.005 with 200 coalition samples ($\approx$ 20% of all possible coalitions). However, when $d=50$, the approximation error remains above 9.0 with 200 samples, stays above 1.0 with $10^{4}$, and only approaches 0.05 near $10^{5}$ coalition samples. These results highlight that as the number of features increases, a substantially larger number of samples is required to obtain reliable estimates—rendering sampling-based methods impractical in high-dimensional settings. This demonstrates the advantage of our exact recursive computation.
@@
Experiment 2: Effectiveness of the Functional-Decomposition-Based Value Function. Recall that PKeX-Shapley employs a value function based on functional decomposition, in contrast to standard expectation-based value functions. It is therefore natural to ask whether this formulation meaningfully captures the importance of feature subsets. To evaluate this, we assess the quality of the resulting explanations in recovering the most informative feature on synthetic datasets, and compare PKeX-Shapley against alternative explanation methods. We generate three regression tasks of $n=1000$ samples in $\mathbb{R}^{50}$, where only the first one-third of features (denoted by $\mathcal{S}$, $\lvert\mathcal{S}\rvert=17$) drive the target, and the remaining 33 features are redundant. The three target functions over $\mathcal{S}$ are: a degree-5 polynomial, a degree-10 polynomial, and a squared-exponential response $y=\exp(\sum_{i\in\mathcal{S}}x_{i}^{2})$. We train a support vector regressor with an RBF kernel on each dataset, and produce explanations using our exact recursive method alongside three baselines: RKHS-SHAP [8], GEMFIX [39], BiSHAP [40], and Sampling SHAP [41], each configured with 500 and 1000 coalition samples. All methods employ a fixed background set of 100 points to estimate their value functions.
Attribution accuracy is computed over 100 independent test instances by selecting the top-17 features returned by each method and measuring the fraction of true active features recovered. Figure 2 shows the average accuracy rate for each method across the three tasks. PKeX-Shapley achieves a competitive or superior performance in all cases, whereas Kernel SHAP, GEMFIX, and Sampling SHAP suffer accuracy degradation as the target function’s complexity increases (most notably for degree-10 polynomial and the exponential model).

### v2

@@ 3.2 A recursive formulation for computing Shapley values
Let $\mathcal{Z}:=\{\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{d}\}$ with $\boldsymbol{z}_{i}=k_{i}(\mathbf{X}_{i},x_{i})$ and $e_{q}(\mathcal{Z})$ the elementary symmetric polynomials (ESPs) of order $q$ over $\mathcal{Z}$, defined recursively as
where $p_{r}(\mathcal{Z})=\sum_{\boldsymbol{z}_{i}\in\mathcal{Z}}\boldsymbol{z}_{i}^{r}$ is the element-wise degree-$r$ power sum. For product-kernel learning methods with coefficients $\boldsymbol{\alpha}$, the Shapley value $\phi_{j}^{\boldsymbol{x}}$ for feature $j$ of instance $\boldsymbol{x}$ can then be expressed as
@@
$\phi_{j}^{\boldsymbol{x}}:=\sum_{\mathcal{S}\subseteq\mathcal{D}\setminus\{j\}}\mu(|\mathcal{S}|)\big(v_{\boldsymbol{x}}(\mathcal{S}\cup\{j\})-v_{\boldsymbol{x}}(\mathcal{S})\big)=\boldsymbol{\alpha}^{\top}\bigg(\big(k_{j}(\mathbf{X}_{j},x_{j})-\boldsymbol{1}\bigg)\bigodot\sum_{q=0}^{d-1}\mu(q)e_{q}\big(\mathcal{Z}_{-j}\big)\bigg).$ (4)
@@
The recursion follows from Newton’s identities for ESPs (Egge, 2019), which due to space constraints, is explained further in Appendix C. Here we provide the key intuition behind equation (4). The multiplicative structure of the product kernel allows us to factorize the marginal contribution $v_{\boldsymbol{x}}(\mathcal{S}\cup\{j\})-v_{\boldsymbol{x}}(\mathcal{S})$ as $\boldsymbol{\alpha}^{\top}((k_{j}(\mathbf{X}_{j},x_{j})-1)\odot k_{S}(\mathbf{X}_{\mathcal{S}},\boldsymbol{x}_{\mathcal{S}}))$. This structure allows us then to push the summation inside the inner product between the coefficients $\boldsymbol{\alpha}$ and the kernel evaluations, and express the total sum over all subsets $\mathcal{S}\subseteq\mathcal{D}\setminus\{j\}$ in terms of weighted ESPs $e_{q}(\mathcal{Z}_{-j})$, which can then be computed recursively in $O(d^{2})$. A similar trick has been used in the context of additive Gaussian processes (Duvenaud et al., 2011), but to our knowledge, it has not been previously leveraged for computing Shapley values in polynomial time. The full algorithm is given in Appendix E, together with a numerically stable modification.
@@
Next, we show the additivity of the explanation for the value function in (3) (see Appendix F for discussion on the additivity of explanations for different values for the null game).
@@
Explaining kernel-based statistical discrepancies
While most explainability research has focused on predictive models [^1] Fleissner et al. (2024) designed an explanation algorithm for kernel-based unsupervised learning., far less attention has been given to statistical inference tasks, where kernel methods are equally powerful. To both demonstrate the flexibility of our framework and advocate for interpretable statistical inference, we show how it applies to the Maximum Mean Discrepancy (MMD) (Gretton et al., 2006) and the Hilbert–Schmidt Independence Criterion (HSIC) (Gretton et al., 2007), two widely used measures of distributional discrepancy and independence.
4.1 Distributing the discrepancy: Explaining the MMD
The Maximum Mean Discrepancy (MMD) quantifies the difference between two probability distributions in terms of their kernel mean embeddings: $\text{MMD}^{2}(\mathbb{P},\mathbb{Q}):=\|\mu_{\mathbb{P}}-\mu_{\mathbb{Q}}\|_{\mathcal{H}_{k}}^{2}$, where $\mathcal{H}_{k}$ is the RKHS associated with the kernel $k$ and $\mu_{\mathbb{P}}:=\int k(\boldsymbol{x},\cdot)\,d\mathbb{P}(\boldsymbol{x})\in\mathcal{H}_{k}$ the kernel mean embedding of $\mathbb{P}$ (Muandet et al., 2017). The embedding $\mu_{\mathbb{Q}}$ is defined analogously. Based on the samples $\{\boldsymbol{x}^{(i)}\}_{i=1}^{n}\overset{i.i.d}{\sim}\mathbb{P}$ and $\{\boldsymbol{z}^{(i)}\}_{i=1}^{m}\overset{i.i.d}{\sim}\mathbb{Q}$, the empirical estimate of MMD, expressed entirely in terms of $k$, is given by $\widehat{\text{MMD}}^{2}(\mathbb{P},\mathbb{Q})=\frac{1}{n(n-1)}\sum_{i\neq j}k(\boldsymbol{x}^{(i)},\boldsymbol{x}^{(j)})+\frac{1}{m(m-1)}\sum_{i\neq j}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})-\frac{2}{nm}\sum_{i,j}k(\boldsymbol{x}^{(i)},\boldsymbol{z}^{(j)}).$
In this section, we propose an attribution method to allocate the overall discrepancy measured by MMD among the involved variables. Such an attribution is useful in many problems, including explaining MMD-based statistics for hypothesis testing (e.g., Gretton et al. (2012)) and determining the contribution of variables to covariance shift (e.g., Zhang et al. (2020)). Similar to the motivation presented in Section 3, as inspired by the multiplicative structure of product kernels, we define the value function of a coalition $\mathcal{S}\subseteq\mathcal{D}$ as the MMD computed only on the kernel components corresponding to $\mathcal{S}$, with the remaining components set to their least informative state (equal to $1$ since we are doing multiplication). This isolates the discrepancy attributable to $\mathcal{S}$ alone.
Definition 5.
For a coalition $\mathcal{S}\subseteq\mathcal{D}$, the functional baseline value function for MMD is defined as $v_{\text{MMD}}(\mathcal{S})=\frac{1}{n(n-1)}\sum_{i\neq j}k_{\mathcal{S}}(\boldsymbol{x}^{(i)}_{\mathcal{S}},\boldsymbol{x}^{(j)}_{\mathcal{S}})+\frac{1}{m(m-1)}\sum_{i\neq j}k_{\mathcal{S}}(\boldsymbol{z}^{(i)}_{\mathcal{S}},\boldsymbol{z}^{(j)}_{\mathcal{S}})-\frac{2}{nm}\sum_{i,j}k_{\mathcal{S}}(\boldsymbol{x}^{(i)}_{\mathcal{S}},\boldsymbol{z}^{(j)}_{\mathcal{S}})$.
@@ 4.1 Distributing the discrepancy: Explaining the MMD
4.2 Distributing the dependence: Explaining the HSIC
The Hilbert-Schmidt Independence Criterion (HSIC) is a kernel-based dependence measure that quantifies the statistical dependence between two random variables. Let $X$ and $Y$ be two random variables with $k(\cdot,\cdot)$ and $l(\cdot,\cdot)$ as reproducing (product) kernels defined on them. Then, $\text{HSIC}(X,Y):=\|\mathcal{C}_{XY}\|^{2}_{\text{HS}}$, where $\mathcal{C}_{XY}$ is the cross-covariance operator and $\|\cdot\|_{\text{HS}}$ is the Hilbert-Schmidt (HS) norm; see, e.g., (Muandet et al., 2017, Sec. 3.6) for technical details. Given a sample $\{(\boldsymbol{x}^{(i)},\boldsymbol{y}^{(i)})\}_{i=1}^{n}\overset{i.i.d}{\sim}\mathbb{P}(X,Y)$, $\text{HSIC}(X,Y)$ can be estimated as $\widehat{\text{HSIC}}(X,Y)=(n-1)^{-2}\mathrm{tr}(\mathbf{K}\mathbf{H}\mathbf{L}\mathbf{H})$ where $\mathbf{K}\in\mathbb{R}^{n\times n}$ is the kernel matrix computed using the kernel $k$, i.e., $\mathbf{K}_{ij}=k(\boldsymbol{x}^{(i)},\boldsymbol{x}^{(j)})$, $\mathbf{L}\in\mathbb{R}^{n\times n}$ is the kernel matrix computed using the kernel $l$, i.e, $\mathbf{L}_{ij}=l(\boldsymbol{y}^{(i)},\boldsymbol{y}^{(j)})$, and $\mathbf{H}=I-\frac{1}{n}\boldsymbol{1}\boldsymbol{1}^{\top}$ is the centering matrix ensuring zero mean in the feature space.
Computing $\text{HSIC}(X,Y)$ gives us the overall dependence between $X$ and $Y$. We now focus on scenarios where $X$ represents a random variable of the feature vector $\boldsymbol{x}$ and $Y$ represents the scalar prediction outcome $y$. Our interest is then to distribute the overall dependence over the features in $\boldsymbol{x}$. This is particularly useful for feature selection and global sensitivity analysis, where the target is usually univariate. Analogous to the predictive function and MMD cases, we define the functional baseline value function of a coalition $\mathcal{S}\subseteq\mathcal{D}$ as the HSIC computed only with the kernel restricted to $\mathcal{S}$, with all other kernel components set to one. This isolates the dependence captured by $\mathcal{S}$ alone:
Definition 7.
For a coalition $\mathcal{S}\subseteq\mathcal{D}$, the functional baseline value function for HSIC is defined as $v_{\text{HSIC}}(\mathcal{S})\;=\;\frac{1}{(n-1)^{2}}\,\mathrm{tr}\!\left(\mathbf{H}\mathbf{L}\mathbf{H}\,\mathbf{K}_{\mathcal{S}}\right),$ where $\mathbf{K}_{\mathcal{S}}=\bigodot_{j\in\mathcal{S}}\mathbf{K}_{j}$ is the Hadamard product of kernel matrices restricted to $\mathcal{S}$.
This construction follows the same principle as before: setting kernel contributions of features outside $\mathcal{S}$ to their least informative value (equal to $1$ as we are doing multiplication). Thus $v_{\text{HSIC}}(\mathcal{S})$ quantifies the dependency between $X_{\mathcal{S}}$ and $Y$, independently of the remaining features. Applying the same trick as in Proposition 6 to the value function $v_{\text{HSIC}}$ yields a similar recursive formulation of Shapley values for HSIC.
@@ 4.2 Distributing the dependence: Explaining the HSIC
$\phi_{j}^{\text{HSIC}}=\frac{1}{(n-1)^{2}}\mathrm{tr}\left(\mathbf{H}\mathbf{L}\mathbf{H}\big((\mathbf{K}_{j}-\boldsymbol{1}\boldsymbol{1}^{\top})\bigodot\sum_{q=0}^{d-1}\mu(q)E_{q}(\mathcal{K}_{-j})\big)\right),{}$
@@
The Shapley values $\phi_{j}^{\text{HSIC}}$ allow us to allocate the overall dependence between $X$ and $Y$ across individual features, identifying the most influential ones for prediction. Our results can be generalized to scenarios when both $X$ and $Y$ are multivariate; see Appendix G for more details. The attribution of overall dependence to the involved multivariate variables has other applications in statistical inference, e.g., kernel-based (conditional) independence testing (Gretton et al., 2007; Albert et al., 2022).
Experiments
@@
We evaluate the effectiveness of PKeX-Shapley for product-kernel methods through a series of experiments conducted on a 24-core machine with 16GB RAM. We further provide the experimental setups, training procedure, and extra experiments in Appendix H.
5.1 Effectiveness of recursion and value function in Local Explanations
@@
Experiment 1: Recursion vs. Regression Formulation. We empirically demonstrate the advantage of PKeX-Shapley, compared to sampling-based approximations such as Kernel SHAP. We generate four synthetic datasets, each with $1000$ samples and $d\in\{10,20,30,50\}$ features. Features are independently drawn from a standard normal distribution, and labels are generated using a linear model with additive Gaussian noise ($\sigma=0.1$). For each dataset, we train a support vector regression model with an RBF kernel and use the trained model to compare explanation methods.
@@
For a fair comparison between the recursive and regression-based methods, we adopt the same functional baseline value function as in (3). Specifically, we first compute the exact Shapley values $\phi_{j}^{\boldsymbol{x}}$ using PKeX-Shapley, and then estimate the Shapley values using a regression-based approach analogous to Kernel SHAP, but modified to employ our value function. The estimator uses the paired coalition sampling scheme of Covert and Lee (2021), with sample sizes ranging from $200$ to $10^{5}$, and computes $\hat{\phi}_{j}^{\boldsymbol{x}}$ by solving the corresponding weighted linear regression problem. To quantify the approximation error introduced by the regression formulation, we report the relative deviation error across $100$ randomly selected instances, defined as $\sum_{j=1}^{d}|\phi_{j}^{\boldsymbol{x}}-\hat{\phi}_{j}^{\boldsymbol{x}}|/|\phi_{j}^{\boldsymbol{x}}|$.
@@
Figure 2 plots the relative error as a function of sample size for each dataset over the 100 selected instances. When $d=10$, the relative error is already around 0.005 with 200 coalition samples ($\approx$ 20% of all possible coalitions). However, when $d=50$, the approximation error remains above 9.0 with 200 samples, stays above 1.0 with $10^{4}$, and only approaches 0.05 near $10^{5}$ coalition samples. These results highlight that as the number of features increases, a substantially larger number of samples is required to obtain reliable estimates—rendering sampling-based methods unreliable in high-dimensional settings. This, in turn, supports the importance of having a fast and exact computation algorithm.
@@
Experiment 2: Effectiveness of the Functional Baseline Value Function. Recall that PKeX-Shapley employs a functional baseline value function, in contrast to standard expectation-based value functions. It is therefore natural to ask whether this formulation meaningfully captures the importance of feature subsets. To evaluate this, we assess the quality of the resulting explanations in recovering the most informative feature on synthetic datasets, and compare PKeX-Shapley against alternative explanation methods. We generate three regression tasks of $n=1000$ samples in $\mathbb{R}^{50}$, where only the first one-third of features (denoted by $\mathcal{S}$, $\lvert\mathcal{S}\rvert=17$) drive the target, and the remaining 33 features are redundant. The three target functions over $\mathcal{S}$ are: a degree-5 polynomial, a degree-10 polynomial, and a squared-exponential response $y=\exp(\sum_{i\in\mathcal{S}}x_{i}^{2})$. We train a support vector regressor with an RBF kernel on each dataset, and produce explanations using our exact recursive method alongside three baselines: RKHS-SHAP (Chau et al., 2022), GEMFIX (Mohammadi et al., 2025a), BiSHAP (Masoomi et al., 2021), and Sampling SHAP (Štrumbelj and Kononenko, 2014), each configured with 500 and 1000 coalition samples. All methods employ a fixed background set of 100 points to estimate their value functions.
Attribution accuracy is computed over 100 independent test instances by selecting the top-17 features returned by each method and measuring the fraction of true active features recovered. Figure 3 shows the average accuracy rate for each method across the three tasks. PKeX-Shapley achieves a competitive or superior performance in all cases, whereas Kernel SHAP, GEMFIX, and Sampling SHAP suffer accuracy degradation as the target function’s complexity increases (most notably for the degree-10 polynomial and the exponential model).